# Landau.bib

### v1

@@ 2.1 Kernel assumptions
We consider the fuzzy Landau equations (1.1) on $\Omega\times\mathbb{R}^{3}$, $\mathbb{R}^{3}\times\mathbb{R}^{3}$ or $\mathbb{R}^{2d}$ for some $d\geq 2$. The spatial kernel $\kappa$ and the interaction kernel
@@
  1. Case 1.
    On the domain $\mathbb{T}^{3}\times\mathbb{R}^{3}$, the kernel $A:\mathbb{R}^{3}\to\mathbb{R}_{+}$ is given by
    $A(v-v_{*})=|v-v_{*}|^{2+\gamma}\quad\text{with}\quad\gamma\in(-3,1].$ (2.1)
    The spatial kernel $k:\mathbb{T}^{3}\to\mathbb{R}_{+}$ is given by
@@
    for some constants ${k_{1},\,k_{2}}>0$ such that $\|\kappa\|_{L^{1}(\mathbb{T}^{3})}=1$.
  2. Case 2.
    On the domain $\mathbb{R}^{3}\times\mathbb{R}^{3}$, the kernel $A:\mathbb{R}^{3}\to\mathbb{R}_{+}$ is given by (2.1). The spatial kernel $k:\mathbb{R}^{3}\to\mathbb{R}_{+}$ satisfies
@@
Remark 2.1 (Dimension discussion).
The cases of kernels satisfying Assumption 2.1 are restricted to dimension three, where the relation between Fisher information and entropy dissipation (see Section 2.4 below) is used to show Theorem 1.1. More precisely, in dimension three, the following cross-product identity holds
$|x|^{2}(y\cdot\Pi_{x^{\perp}}y)=|x\times y|^{2}\quad\forall\,x,\,y\in\mathbb{R}^{3}.$ (2.5)
See for example [carrillo2024landau, Lemma 42] for a proof.
On the other hand, the variational characterisation theorem (1.1) holds also for a pair of kernels $(\bar{A},\kappa)$ defined below for all dimension $\mathbb{R}^{2d}$, $d\geq 2$.
The kernels ${\bar{A}}:\mathbb{R}^{d}\to\mathbb{R}_{+}$ are measurable, continuous, and symmetric in the sense that
$\bar{A}(v-v_{*})=\bar{A}(v_{*}-v)\quad\forall v,\,v_{*}\in\mathbb{R}^{d}.$
Moreover, for $\gamma\in(-\infty,3]$, the following bounds hold
$C_{\bar{A}}^{-1}\langle v-v_{*}\rangle^{\gamma}\leq\bar{A}(v-v_{*})\leq C_{\bar{A}}\langle v-v_{*}\rangle^{\gamma}$
for some constant $C_{\bar{A}}>0$. The spatial kernel $\kappa$ is defined as in (2.2). The pair of kernels $({\bar{A}},\kappa)$ were considered for the fuzzy Boltzmann equation in [EH24]. Theorem 1.1 holds for such kernel $({\bar{A}},\kappa)$ in a slightly different framework, where we consider an unweighted gradient
~$\tilde{\nabla}_{L}f=\Pi_{(v-v_{*})^{\perp}}\big{(}\nabla_{v}f-\nabla_{v_{*}}f_{*}\big{)}.$
Details can be found in Section 4.
@@
In Assumption 2.1, spatial kernels $\kappa$ has positive upper and lower bounds
$C^{-1}\leq\kappa\leq C,\quad\text{for some constant }C>0.$
Notice that $\kappa=1$ is one example that satisfies assumptions (2.3)-(2.4) in Case 2.4.
Let $M(z)=C\exp\big{(}-\langle z\rangle\big{)}$, where we choose the constant $C>0$ such that $\|M\|_{L^{1}(\mathbb{R}^{3})}=1$. We define the renormalised sequence of molliffiers with parameter $\beta\in(0,1)$
$M_{\beta}(z)=C\beta^{-3}\exp(-\langle{z}/{\beta}\rangle).$ (2.8)
@@
Notice that the uniform bounds (2.9) also hold for interaction kernels defined in Remark 2.1
${\bar{A}}*_{v}M_{\beta}\leq C{\bar{A}}$ $\text{and}\quad{\bar{A}}^{-1}*_{v}M_{\beta}\leq C{\bar{A}}^{-1}.$ (2.10)
The proof of (2.10) can be found in [EH24, Lemma 2.1], which is a direct consequence of Peetre’s inequality
$\frac{\langle x\rangle^{p}}{\langle y\rangle^{p}}\leq 2^{|p|/2}\langle x-y\rangle^{|p|}\quad\forall\,x,\,y\in\mathbb{R}^{3},\,\,p\in\mathbb{R},$
@@
where $C>0$ depends only on $|p|$ and $\|\langle\cdot\rangle^{|p|}M(\cdot)\|_{L^{1}(\mathbb{R}^{3})}$.
@@ 2.2 Notations
For simplification, we write the domain as $\mathbb{R}^{3}\times\mathbb{R}^{3}$, and unless otherwise stated, it refers to both $\mathbb{R}^{3}\times\mathbb{R}^{3}$ and $\mathbb{T}^{3}\times\mathbb{R}^{3}$.
@@
$L^{1}_{a,b}(\mathbb{R}^{3}\times\mathbb{R}^{3})$ =$=\{f\in L^{1}(\mathbb{R}^{3}\times\mathbb{R}^{3})\mid\|f\|_{L^{1}_{a,b}(\mathbb{R}^{3}\times\mathbb{R}^{3})}<+\infty\},$
@@
$\|f\|_{L^{1}_{a,b}(\mathbb{R}^{3}\times\mathbb{R}^{3})}\mathop{=}\limits^{\textrm{def}}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\big{(}\langle x\rangle^{a}+\langle v\rangle^{b}\big{)}f(x,v){\,\rm d}x{\,\rm d}v,$
where we recall that $\langle z\rangle:=\sqrt{1+|z|^{2}}$ denotes the Japanese jacket. Let $p,\,q\in[1,\infty)$. The functional space $L^{p}_{v}L^{q}_{x}(\mathbb{R}^{3}\times\mathbb{R}^{3})$ consists of all the measure functions $f$ such that
$\|f\|_{L^{p}_{v}L^{q}_{x}}\mathop{=}\limits^{\textrm{def}}\Big{(}\int_{\mathbb{R}^{3}}\Big{(}\int_{\mathbb{R}^{3}}|f(x,v)|^{q}{\,\rm d}x\Big{)}^{\frac{p}{q}}{\,\rm d}v\Big{)}^{\frac{1}{p}}<+\infty.$
We define $\mathcal{P}(\mathbb{R}^{3}\times\mathbb{R}^{3})$ as the space of Borel probability measures on $\mathbb{R}^{3}\times\mathbb{R}^{3}$, which is endowed with the weak topology induced by duality with bounded continuous functions in $C_{b}(\mathbb{R}^{3}\times\mathbb{R}^{3})$.
For any $a,\,b\in\mathbb{R}$, we define the following subspace of $\mathcal{P}(\mathbb{R}^{3}\times\mathbb{R}^{3})$
$\mathcal{P}_{a,b}(\mathbb{R}^{3}\times\mathbb{R}^{3})=\{\mu\in\mathcal{P}(\mathbb{R}^{3}\times\mathbb{R}^{3})\mid\mathcal{E}_{a,b}(\mu)<+\infty\},$
@@
$\mathcal{E}_{a,b}(\mu)\mathop{=}\limits^{\textrm{def}}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}(\langle x\rangle^{a}+\langle v\rangle^{b}){\,\rm d}\mu.$
For functions $f:\mathbb{R}^{3}\times\mathbb{R}^{3}\to\mathbb{R}$, we also write
$\mathcal{E}_{a,b}(f)=\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\big{(}\langle x\rangle^{a}+\langle v\rangle^{b}\big{)}f(x,v){\,\rm d}x{\,\rm d}v.$
Next we recall the Boltzmann entropy functional for measure $\mu\in\mathcal{P}(\mathbb{R}^{3}\times\mathbb{R}^{3})$. If $\mu=f\mathcal{L}$ is absolutely continuous with respect to Lebesgue measure $\mathcal{L}$ with density $f$, we define
$\mathcal{H}(\mu)=\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f\log f{\,\rm d}x{\,\rm d}v.$
Otherwise, we define $\mathcal{H}(\mu)=+\infty$. In the case of $\mu=f\mathcal{L}$, we also write $\mathcal{H}(f)$. An important property will be used in the subsequent analysis is that the entropy $\mathcal{H}(\mu)$ is weakly lower semi-continuous in the sense that if $\mu_{n}\to\mu$ weakly in $\mathcal{P}(\mathbb{R}^{3}\times\mathbb{R}^{3})$ and $\mathcal{E}_{2,2}(\mu_{n})\to\mathcal{E}_{2,2}(\mu)$ as $n\to\infty$, then
@@ 2.3 Assumptions on curves
$(\partial_{t}+v\cdot\nabla_{x})f=\frac{1}{2}\tilde{\nabla}\cdot\big{(}\kappa ff_{*}\tilde{\nabla}\log f\big{)},$ (2.12)
where we call that the gradient operator $\tilde{\nabla}$ is defined in (1.11) and the divergence operator $\tilde{\nabla}\cdot$ is defined via the integration by parts formula (1.12)
$\int_{\mathbb{R}^{12}}\tilde{\nabla}g\cdot B{\,\rm d}\eta=-\int_{\mathbb{R}^{6}}g\tilde{\nabla}\cdot B{\,\rm d}v{\,\rm d}x$
for all $B(x,x_{*},v,v_{*}):(\mathbb{R}^{3}\times\mathbb{R}^{3})^{2}\to\mathbb{R}^{3}$ and $g:\mathbb{R}^{3}\times\mathbb{R}^{3}\to\mathbb{R}$. We follow [Vil98b] to define a $\mathcal{H}$-solution for fuzzy Landau equation (1.1).
@@
Let $\gamma\in(-3,1]$ and $A(v-v_{*})=|v-v_{*}|^{2+\gamma}$ is defined as in (2.1). Let
$f\in C([0,T];L^{1}(\mathbb{R}^{3}\times\mathbb{R}^{3}))\cap L^{1}([0,T];L^{1}_{2,2+\gamma_{+}}(\mathbb{R}^{3}\times\mathbb{R}^{3})),$
where $\gamma_{+}=\max\{\gamma,0\}$. If $\gamma\in(-3,-2)$, then additionally we assume
$f\in L^{1}([0,T];L^{p}_{v}L^{1}_{x})$ (2.13)
for some $\frac{3}{3-|2+\gamma|}<p$. We say $f$ is a $\mathcal{H}$-solution of (1.1) if
$f_{t}\geq 0\quad\text{and}\quad\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f_{t}(x,v){\,\rm d}x{\,\rm d}v=1\quad\forall\,t\in[0,T];$
@@
$\int_{0}^{T}\mathcal{D}(f_{t}){\,\rm d}t=\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{12}}\kappa ff_{*}|\tilde{\nabla}\log f|^{2}{\,\rm d}\eta{\,\rm d}t<+\infty;$
@@
$\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f_{0}\varphi(0){\,\rm d}x{\,\rm d}v+\int_{0}^{T}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(\partial_{t}\varphi+v\cdot\nabla_{x}\varphi){\,\rm d}x{\,\rm d}v{\,\rm d}t$ (2.14)
$=-\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{12}}\kappa ff_{*}\tilde{\nabla}\varphi\cdot\tilde{\nabla}\log f{\,\rm d}\eta{\,\rm d}t$
for all $\varphi\in C^{\infty}_{c}([0,T)\times\mathbb{R}^{3}\times\mathbb{R}^{3})$.
Notice that, on the domain $\mathbb{T}^{3}\times\mathbb{R}^{3}$, we do not need extra moments assumption on $x$.
@@
  • The right-hand side of (2.14) is well-defined. Indeed, by the Cauchy–Schwarz inequality,
    $\Big{|}\int_{0}^{T}\int_{\mathbb{R}^{12}}\kappa ff_{*}\tilde{\nabla}\varphi\cdot\tilde{\nabla}\log f{\,\rm d}\eta{\,\rm d}t\Big{|}$
    $\leq\Big{|}\int_{0}^{T}\mathcal{D}(f){\,\rm d}t\Big{|}^{\frac{1}{2}}\Big{|}\int_{0}^{T}\int_{\mathbb{R}^{12}}\kappa ff_{*}|\tilde{\nabla}\varphi|^{2}{\,\rm d}\eta{\,\rm d}t\Big{|}^{\frac{1}{2}}.$
@@
    $\kappa ff_{*}|\tilde{\nabla}\varphi|^{2}\leq C(\varphi,\gamma)ff_{*}(|v|^{2+\gamma}+|v_{*}|^{2+\gamma})\in L^{1}([0,T]\times\mathbb{R}^{12}).$
    In the case of very soft potential $\gamma\in(-3,-2)$, we apply Corollary 2.8 below to derive
    $\Big{|}\int_{0}^{T}\int_{\mathbb{R}^{12}}\kappa ff_{*}|\tilde{\nabla}\varphi|^{2}{\,\rm d}\eta{\,\rm d}t\Big{|}$
    $\leq$ ${}C(\varphi)\int_{0}^{T}\int_{\mathbb{R}^{12}}ff_{*}|v-v_{*}|^{2+\gamma}{\,\rm d}\eta{\,\rm d}t$
    $\lesssim$ ${}\|\langle v\rangle^{|2+\gamma|}f\|_{L^{1}([0,T];L^{p}_{v}L^{1})}.$
  • The assumption (2.13) is unnecessary in the homogeneous Landau cases in [Vil98b]. In the homogeneous case with $\psi=\psi(v)\in C^{\infty}_{c}(\mathbb{R}^{3})$, we have
    ~$\tilde{\nabla}_{\sf homo}\psi(v)=|v-v_{*}|^{1+\frac{\gamma}{2}}\Pi_{(v-v_{*})^{\perp}}\big{(}\nabla_{v}\psi(v)-\nabla_{v_{*}}\psi(v_{*})\big{)},$
@@
    |~$|\tilde{\nabla}_{\sf homo}\psi|$ $\leq|v-v_{*}|^{1+\frac{\gamma}{2}}|\nabla_{v}\psi(v)-\nabla_{v_{*}}\psi(v_{*})|$
    $\leq|v-v_{*}|^{2+\frac{\gamma}{2}}\|\psi\|_{\operatorname{Lip}(\mathbb{R}^{3})},$
    where $2+\frac{\gamma}{2}\geq\frac{1}{2}$. However, this is not the case for fuzzy Landau equations, since we also need to take into account the distance between $x$ and $x_{*}$ in
@@
    We note that $f\in L^{1}_{2,2}(\mathbb{R}^{3}\times\mathbb{R}^{3})$ ensures $\mathcal{H}(f_{T})>-\infty$.
@@
  1. Case 1.
    For $(f_{t}\mathcal{L})_{t\in[0,T]}\subset\mathcal{P}_{0,2+|\gamma|}(\mathbb{T}^{3}\times\mathbb{R}^{3})$, we assume
    • in the soft potential case $\gamma\in(-3,0)$
      $f\in L^{\infty}([0,T];L^{\infty}_{0,2+|\gamma|}(\mathbb{T}^{3}\times\mathbb{R}^{3}))\quad\text{and}\quad\langle v\rangle^{2+|\gamma|}f_{t}\in L^{1}([0,T];L^{p}_{v}L^{1}_{x}),$
    • in the Maxwellian potential case $\gamma=0$
      $f\in L^{\infty}([0,T];L^{1}_{0,2}(\mathbb{T}^{3}\times\mathbb{R}^{3})),$
    • in the hard potential case $\gamma\in(0,1]$
      $f\in L^{\infty}([0,T];L^{1}_{0,2+\gamma}(\mathbb{T}^{3}\times\mathbb{R}^{3}))\quad\text{and}\quad f\in L^{\infty}([0,T];L^{p}_{v}L^{1}_{x})$
    for some $p\in\big{(}\frac{3}{3-|\gamma|},+\infty\big{)}$.
  2. Case 2.
    For $(f_{t}\mathcal{L})_{t\in[0,T]}\subset\mathcal{P}_{2,2+|\gamma|}(\mathbb{R}^{3}\times\mathbb{R}^{3})$, we assume $f$ satisfies the assumptions in Case 1 for all $\gamma\in(-3,1]$. In addition, we always assume
    $f\in L^{1}([0,T];L^{1}_{2,0}(\mathbb{R}^{3}\times\mathbb{R}^{3})).$

### v2

@@ 2 Preliminary
Throughout this article, for simplification, we write the domain as $\mathbb{R}^{2d}$, and unless otherwise stated, it refers to both $\mathbb{R}^{2d}$ and $\mathbb{T}^{d}\times\mathbb{R}^{d}$.
@@ 2.1 Kernel assumptions
The spatial kernel $\kappa$ and the interaction kernel
@@
The kernel $A:\mathbb{R}^{d}\to\mathbb{R}_{+}$ is given by
$A(v-v_{*})=|v-v_{*}|^{2+\gamma}\quad\text{with}\quad\gamma\in(-\gamma_{d},1].$ (2.1)
The spatial kernel $k:\mathbb{R}^{d}\to\mathbb{R}_{+}$ is given by
@@
Remark 2.2.
In Assumption 2.1, we can also consider the spatial kernel $k:\mathbb{R}^{d}\to\mathbb{R}_{+}$ with positive lower and upper bounds
@@
for some constant $C>0$. Notice that $\kappa=1$ is included in this case.
Remark 2.3.
In Assumption 2.1, in the hard potential cases, one can also take the kernel $A$ of the following form
$A(v-v_{*})=\overline{A}(v-v_{*})|v-v_{*}|^{2},\quad\forall\,v,\,v_{*}\in\mathbb{R}^{d},$ (2.5)
where $\overline{A}(v-v_{*})=\overline{A}(v_{*}-v)$ and the following bounds hold
$C^{-1}\langle v-v_{*}\rangle^{\gamma+2}\leq\overline{A}(v-v_{*})\leq C\langle v-v_{*}\rangle^{\gamma+2},\quad\gamma\in(-\infty,1]$ (2.6)
@@
The spatial kernel $\kappa$ is defined as in (2.2). The pair of kernels $({\overline{A}},\kappa)$ were considered for the fuzzy Boltzmann equation in [EH25]. The fuzzy Landau equation (1.1) with kernel (2.5) is the grazing limit of the fuzzy Boltzmann equation in [EH25] with non-cutoff angular kernels, more details can be found in [DGH25].
@@
Let $M(z)=C\exp\big(-\langle z\rangle\big)$, where we choose the constant $C>0$ such that $\|M\|_{L^{1}(\mathbb{R}^{d})}=1$. We define the renormalised sequence of mollifiers with parameter $\beta\in(0,1)$
$M_{\beta}(z)=C\beta^{-d}\exp(-\langle{z}/{\beta}\rangle).$ (2.9)
@@
The proof of (2.10) can be found in [EH25, Lemma 2.1], which is a direct consequence of Peetre’s inequality
$\frac{\langle x\rangle^{p}}{\langle y\rangle^{p}}\leq 2^{|p|/2}\langle x-y\rangle^{|p|}\quad\forall\,x,\,y\in\mathbb{R}^{d},\,\,p\in\mathbb{R},$
@@
where $C>0$ depends only on $|p|$ and $\|\langle\cdot\rangle^{|p|}M(\cdot)\|_{L^{1}(\mathbb{R}^{d})}$.
@@ 2.2 Notations
For simplification, we write the domain as $\mathbb{R}^{2d}$, and unless otherwise stated, it refers to both $\mathbb{R}^{2d}$ and $\mathbb{T}^{d}\times\mathbb{R}^{d}$.
@@
$L^{1}_{a,b}(\mathbb{R}^{2d})$ $=\{f\in L^{1}(\mathbb{R}^{2d})\mid\|f\|_{L^{1}_{a,b}(\mathbb{R}^{2d})}<+\infty\},$
@@
$\|f\|_{L^{1}_{a,b}(\mathbb{R}^{2d})}\mathop{=}\limits^{\textrm{def}}\int_{\mathbb{R}^{2d}}\big(\langle x\rangle^{a}+\langle v\rangle^{b}\big)f(x,v){\,\rm d}x{\,\rm d}v,$
where we recall that $\langle z\rangle:=\sqrt{1+|z|^{2}}$ denotes the Japanese jacket. Let $p,\,q\in[1,\infty)$. The functional space $L^{p}_{v}L^{q}_{x}(\mathbb{R}^{2d})$ consists of all the measure functions $f$ such that
$\|f\|_{L^{p}_{v}L^{q}_{x}}\mathop{=}\limits^{\textrm{def}}\Big(\int_{\mathbb{R}^{d}}\Big(\int_{\mathbb{R}^{d}}|f(x,v)|^{q}{\,\rm d}x\Big)^{\frac{p}{q}}{\,\rm d}v\Big)^{\frac{1}{p}}<+\infty.$
Similarly, we can also define the functional spaces $L^{q}_{x}L^{p}_{v}(\mathbb{R}^{2d})$.
We define $\mathcal{P}(\mathbb{R}^{2d})$ as the space of Borel probability measures on $\mathbb{R}^{2d}$, which is endowed with the weak topology induced by duality with bounded continuous functions in $C_{b}(\mathbb{R}^{2d})$.
For any $a,\,b\in\mathbb{R}$, we define the following subspace of $\mathcal{P}(\mathbb{R}^{2d})$
$\mathcal{P}_{a,b}(\mathbb{R}^{2d})=\{\mu\in\mathcal{P}(\mathbb{R}^{2d})\mid\mathcal{E}_{a,b}(\mu)<+\infty\},$
@@
$\mathcal{E}_{a,b}(\mu)\mathop{=}\limits^{\textrm{def}}\int_{\mathbb{R}^{2d}}(\langle x\rangle^{a}+\langle v\rangle^{b}){\,\rm d}\mu.$
For functions $f:\mathbb{R}^{2d}\to\mathbb{R}$, we also write
$\mathcal{E}_{a,b}(f)=\int_{\mathbb{R}^{2d}}\big(\langle x\rangle^{a}+\langle v\rangle^{b}\big)f(x,v){\,\rm d}x{\,\rm d}v.$
Next we recall the Boltzmann entropy functional for measure $\mu\in\mathcal{P}(\mathbb{R}^{2d})$. If $\mu=f\mathcal{L}$ is absolutely continuous with respect to Lebesgue measure $\mathcal{L}$ with density $f$, we define
$\mathcal{H}(\mu)=\int_{\mathbb{R}^{2d}}f\log f{\,\rm d}x{\,\rm d}v.$
Otherwise, we define $\mathcal{H}(\mu)=+\infty$. In the case of $\mu=f\mathcal{L}$, we also write $\mathcal{H}(f)$. An important property will be used in the subsequent analysis is that the entropy $\mathcal{H}(\mu)$ is weakly lower semi-continuous in the sense that if $\mu_{n}\to\mu$ weakly in $\mathcal{P}(\mathbb{R}^{2d})$ and $\mathcal{E}_{2,2}(\mu_{n})\to\mathcal{E}_{2,2}(\mu)$ as $n\to\infty$, then
@@ 2.3 Assumptions on curves
$(\partial_{t}+v\cdot\nabla_{x})f=\frac{1}{2}\widetilde{\nabla}\cdot\big(\kappa ff_{*}\widetilde{\nabla}\log f\big),$ (2.12)
where we call that the gradient operator $\widetilde{\nabla}$ is defined in (1.11) and the divergence operator $\widetilde{\nabla}\cdot$ is defined via the integration by parts formula (1.12)
$\int_{\mathbb{R}^{4d}}\widetilde{\nabla}g\cdot B{\,\rm d}\eta=-\int_{\mathbb{R}^{2d}}g\widetilde{\nabla}\cdot B{\,\rm d}v{\,\rm d}x$
for all $B(x,x_{*},v,v_{*}):\mathbb{R}^{4d}\to\mathbb{R}^{d}$ and $g:\mathbb{R}^{2d}\to\mathbb{R}$. We follow [Vil98a] to define a $\mathcal{H}$-solution for fuzzy Landau equation (1.1).
@@
Let $\gamma\in(-\gamma_{d},1]$ and $A(v-v_{*})=|v-v_{*}|^{2+\gamma}$ is defined as in (2.1). Let
$f\in C([0,T];\mathscr{D}^{\prime}(\mathbb{R}^{2d}))\cap L^{1}([0,T];L^{1}_{2,2+\gamma_{+}}(\mathbb{R}^{2d})),$
where $\gamma_{+}=\max\{\gamma,0\}$. We say $f$ is a $\mathcal{H}$-solution of (1.1) if
$f_{t}\geq 0\quad\text{and}\quad\int_{\mathbb{R}^{2d}}f_{t}(x,v){\,\rm d}x{\,\rm d}v=1\quad\forall\,t\in[0,T];$
@@
$\int_{0}^{T}\mathcal{D}(f_{t}){\,\rm d}t=\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa ff_{*}|\widetilde{\nabla}\log f|^{2}{\,\rm d}\eta{\,\rm d}t<+\infty;$
@@
$\int_{\mathbb{R}^{2d}}f_{0}\varphi(0){\,\rm d}x{\,\rm d}v+\int_{0}^{T}\int_{\mathbb{R}^{2d}}f(\partial_{t}\varphi+v\cdot\nabla_{x}\varphi){\,\rm d}x{\,\rm d}v{\,\rm d}t$ (2.13)
$=-\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa ff_{*}\widetilde{\nabla}\varphi\cdot\widetilde{\nabla}\log f{\,\rm d}\eta{\,\rm d}t$
for all $\varphi\in C^{\infty}_{c}([0,T)\times\mathbb{R}^{2d})$.
Notice that, on the domain $\mathbb{T}^{d}\times\mathbb{R}^{d}$, we do not need an extra moment assumption on $x$.
@@
  • The right-hand side of (2.13) is well-defined. Indeed, by the Cauchy–Schwartz inequality,
    $\Big|\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa ff_{*}\widetilde{\nabla}\varphi\cdot\widetilde{\nabla}\log f{\,\rm d}\eta{\,\rm d}t\Big|$
    $\leq$ ${}\Big|\int_{0}^{T}2\mathcal{D}(f){\,\rm d}t\Big|^{\frac{1}{2}}\Big|\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa ff_{*}|\widetilde{\nabla}\varphi|^{2}{\,\rm d}\eta{\,\rm d}t\Big|^{\frac{1}{2}}.$
@@
    $\kappa ff_{*}|\widetilde{\nabla}\varphi|^{2}\leq C(\varphi,\gamma)ff_{*}(|v|^{2+\gamma}+|v_{*}|^{2+\gamma})\in L^{1}([0,T]\times\mathbb{R}^{4d}).$
    In the case of very soft potential $\gamma\in(-\gamma_{d},-2)$, we assume that $\operatorname{supp}(\varphi)\subset\mathbb{R}^{d}_{x}\times B^{R}_{v}$. By using Hölder and Young’s inequalities, we have
    $\Big|\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa ff_{*}|\widetilde{\nabla}\varphi|^{2}{\,\rm d}\eta{\,\rm d}t\Big|$ (2.14)
    $\leq$ ${}C(\varphi)\int_{0}^{T}\int_{\mathbb{R}^{2d}\times(B^{R})^{2}}ff_{*}|v-v_{*}|^{2+\gamma}{\,\rm d}\eta{\,\rm d}t$
    $\lesssim$ ${}\|f\|_{L^{\infty}([0,T];L^{1}(\mathbb{R}^{2d})}\|f\|_{L^{1}([0,T];L^{1}_{x}L^{p}(B^{R}_{v})}\||v|^{\gamma+2}\|_{L^{p^{\prime}}(B^{2R})}.$
    where $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. The Fisher information estimate in Lemma 2.9 and the Sobolev embedding ensure that $f\in L^{1}([0,T];L^{1}_{x}(L^{k}_{v})_{\operatorname{loc}}$, where $k=\frac{d}{d-2}$ for $d\geq 3$, and $k\in(1,+\infty)$ for $d=2$. We choose $(p,p^{\prime})=(k,d/2)$ in (2.14), then $|\gamma+2|p^{\prime}<d$ ensures that $|v|^{\gamma+2}\in L^{p^{\prime}}_{\operatorname{loc}}(\mathbb{R}^{d})$. The detailed discussion can be found in Part II.
  • In the homogeneous Landau cases in [Vil98a], one can justify the $\mathcal{H}$-solutions in a different way. In the homogeneous case with $\psi=\psi(v)\in C^{\infty}_{c}(\mathbb{R}^{d})$, we have
    $\widetilde{\nabla}_{\sf homo}\psi(v)=|v-v_{*}|^{1+\frac{\gamma}{2}}\Pi_{(v-v_{*})^{\perp}}\big(\nabla_{v}\psi(v)-\nabla_{v_{*}}\psi(v_{*})\big),$
@@
    $|\widetilde{\nabla}_{\sf homo}\psi|$ $\leq|v-v_{*}|^{1+\frac{\gamma}{2}}|\nabla_{v}\psi(v)-\nabla_{v_{*}}\psi(v_{*})|$
    $\leq|v-v_{*}|^{2+\frac{\gamma}{2}}\|\psi\|_{\operatorname{Lip}(\mathbb{R}^{d})},$
    where $2+\frac{\gamma}{2}>0$. However, this is not the case for fuzzy Landau equations, since we also need to take into account the distance between $x$ and $x_{*}$ in
@@
    We note that $f\in L^{1}_{2,2}(\mathbb{R}^{2d})$ ensures $\mathcal{H}(f_{T})>$-$\infty$.
@@
For $(f_{t}\mathcal{L})_{t\in[0,T]}\subset\mathcal{P}_{0,2+|\gamma|}(\mathbb{R}^{2d})$, we assume
$f\in L^{1}\big([0,T];L^{\infty}_{2,0}(\mathbb{R}^{2d})\big)\cap L^{\infty}\big([0,T];L^{\infty}_{0,2+|\gamma|}(\mathbb{R}^{2d})\big).$
In addition, in the soft potential case $\gamma\in(-\gamma_{d},0)$, we assume
$\langle v\rangle^{2+\gamma}f\in L^{\infty}([0,T];L^{p}_{v}L^{1}_{x});$
In the hard potential case $\gamma\in(0,1]$, we assume
$f\in L^{\infty}([0,T];L^{p}_{v}L^{1}_{x})$
for some $p\in\big(\frac{d}{d-|\gamma|},+\infty\big)$.